Assume that adults have IQ scores that are normally distributed with a mean of 101.4 and a standard deviation 20.9. Find the first quartile Upper Q 1​, which is the IQ score separating the bottom​ 25% from the top​ 75%. (Hint: Draw a​ graph.) The first quartile is what. ​(Type an integer or decimal rounded to one decimal place as​ needed.)
Answer #1

SOLUTION:

Using standard normal table,

P(Z < z) = 25%

= P(Z < z) = 0.25  

= P(Z < -0.67) = 0.25

z = -0.67 Using standard normal z table,

Using z-score formula  

x= z * +

x= -0.67 *20.9+101.4

x= 87.4